. In a closed economy, GDP is $1000, government purchases are $200, and consumption is $700. If the government has a budget surplus of $25, what are investment, taxes, private saving, and national saving? Explain with equations. [4 marks]

1
Expert's answer
2021-06-16T09:22:45-0400

In Keynesian macroeconomics when the price level in the economy is assumed to be fixed then planned aggregate expenditure in the economy determines the real GDP level or output level of the economy. The gross domestic product(GDP) refers to the market value of all the final commodities(goods and services) produced inside the domestic boundary of a nation during a particular period of time usually in a year or quarter.

When the value of GDP is calculated at some base year prices then the value of GDP is known as real GDP. The real GDP is the inflation-adjusted or price-adjusted measure of GDP. In Keynesian macroeconomics, the goods market of the economy reaches its equilibrium position when aggregate expenditure in the economy becomes equal to the real GDP level or output level of the economy.


In a closed economy, the aggregate expenditure is equal to the sum of consumption expenditure, government expenditure, and investment expenditure. As we know in the economy given in the question the GDP level(Y) is equal to $100, government purchases(G) is equal to $200 and the consumption expenditure(C) in the economy is equal to $700 so the level of investment expenditure in this economy can be calculated as follows

"Y=C+I+G\\\$1000=\$700+I+\$200\\
I=\$100"


So in this economy, the investment level is equal to $100.

The national saving refers to the portion of the total GDP that is not used by the households in the economy in the form of consumption expenditure or used by the government in the form of government purchases. The national savings level in the economy is calculated as follows


"S=Y-C-G\\S=\$1000-\$700-\$200\\S=\$100"


So in this economy, the saving level is also equal to $100 so we can say that in a closed economy the investment is equal to the national saving.

The budget surplus refers to the excess of government taxes over its expenditure and the budget surplus is also known as the public saving. As we know the budget surplus in the economy is equal to $25 and government purchases are equal to $100 so the taxes in the economy can be calculated as follows


"Budget\space surplus=Taxes\space -government\space expenditure\\\$25=Taxes-\$100\\Taxes=\$125"


In this economy, taxes is equal to $125.

The national saving is always equal to the sum of public saving and private saving as we know that in this economy the public saving is equal to $25 and national saving is equal to $100 so we can calculate the level of private saving in the economy as follow


National saving=public + private saving

"\$100=\$25+private\space saving\\private\space saving=\$75"


In this economy the private saving is equal to $75.